Grade 12Mechanics

A satellite of mass m, initially at rest on the earth, is launched into a circular orbit at a height equal to the radius of the earth. The minimum energy required is

Tanya
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
As we probably am aware, Gravitational potential 
vitality = - GM/r 
what's more, orbital speed, v0 = 
vGM/R+h 
Ef = ½ mv20 
– GMm/3R = ½ m GM/3R – GMm/3R 
= GMm/3R (1/2-1) = - GMm/6R 
Ei = - GM/R + K 
Ei = Ef 
Along these lines 
least required vitality, K = 5GMm/6R